### v3

# ATM: Improving Model Merging by Alternating Tuning and Merging

Abstract
Model merging has recently emerged as a cost-efficient paradigm for multi-task learning. Among current approaches, task arithmetic (Ilharco et al., 2022) stands out for its simplicity and effectiveness. In this paper, we motivate the effectiveness of task vectors by linking them to multi-task gradients. We show that in a single-epoch scenario, if the optimization is performed via gradient descent, task vectors are after one step mathematically equivalent to the gradients obtained via gradient descent in a multi-task setting, and still approximate these gradients in subsequent epochs. Furthermore, we show that the effectiveness of task vectors is largely driven by the first epoch’s gradient. Given this parallel between task vectors and gradients, we propose viewing model merging as a single step in an iterative process that Alternates between Tuning and Merging (ATM). We then propose two ways to utilize ATM. The first is to replace multi-task learning with ATM in scenarios where data sharing is prohibited (e.g. federated learning). The second is to improve the outcome of any model merging algorithm by applying a few post-hoc iterations of ATM on a small validation dataset, which is commonly available for hyperparameter tuning. Finally, we provide both empirical and theoretical support for the effectiveness of ATM, demonstrating that it minimizes an upper bound on the loss obtained by jointly finetuning all tasks.
1 Introduction
The pretrain-and-finetune paradigm has become the standard for many deep learning tasks, where a model pretrained on large-scale, unlabeled data is adapted to a specific downstream task with minimal tuning. However, when working with multiple tasks, a major drawback is the need to store separate finetuned models at inference. Model merging addresses this challenge by combining task-specific models into a single model capable of handling all tasks. This significantly reduces storage costs, as the unified model’s size remains comparable to that of a single-task model, regardless of the number of tasks. Among numerous model merging frameworks, task arithmetic (Ilharco et al., 2022) stands out for its simplicity and effectiveness. Given a pretrained model $\theta_{0}$ and a model $\theta_{i}$ finetuned on task $t_{i}$, the task vector $\tau_{i}=\theta_{i}-\theta_{0}$ is defined as the difference between the finetuned and pretrained weights. For multi-task learning with $n$ tasks, task arithmetic sums the $n$ task vectors, scales the sum with a coefficient $\alpha$, and adds the resulting vector back to the pretrained model.
[…]
Figure 1: The ATM method, illustrated up to iteration $h$ with two tasks (A and B). In each iteration, the Tune step finetunes the pretrained model $\theta_{\text{base}}^{(0)}$ separately on both tasks, and the Merge step aggregates the task vectors and applies the resulting multi-task vector to the base model. This process repeats, with each iteration using the updated model as the new base, continuing until a stopping condition is met.
In this paper, we explain the effectiveness of task arithmetic by linking task vectors to the gradients of the task losses. We start by noticing: when a model is finetuned for a single epoch using gradient descent (GD), the corresponding task vector is the additive inverse of the loss gradient, scaled by the learning rate. Similarly, the multi-task vector, obtained by summing individual task vectors, is equivalent to the additive inverse of the average loss gradient. Thus, task addition is analogous to performing a GD step on the sum of the average task losses. When finetuning spans multiple epochs, this equality becomes an approximation, with an error dependent on the learning rate.
While the single-epoch assumption is not always exact, we confirm that the first epoch contributes most to the gradient norm. Even when it does not, subsequent gradients tend to align with the first, confirming that the initial direction predominantly dictates the task vector’s effectiveness.
In this view, aggregation and merging in task arithmetic correspond to a noisy GD step when finetuning jointly on all tasks, using the sum of the average losses as the objective. In practice, this implies that the one-step nature of these techniques likely leads to overshooting the multi-task optimum, as they effectively perform GD on a multi-task dataset with a single, noisy step. The scaling factors optimized via the validation set essentially act as the learning rates in this process.
Building on these insights, we propose viewing task arithmetic as a single step in an iterative framework and introduce Alternating Tuning and Merging (ATM) – a generic framework that iteratively alternates between finetuning and merging. We propose applying ATM in two ways. The first is PA-ATM (Privacy-Aware ATM), which replaces multi-task learning in scenarios where multi-task data sharing is forbidden (e.g. federated learning). The second is PH-ATM (Post-Hoc ATM), which enhances any multi-task model by applying a few post-hoc iterations of ATM on a small set of validation data, which is commonly available and used for hyperparameter tuning. The flexibility of ATM enables the integration of any conflict-resolution method in the task arithmetic literature for further gains.
Finally, we rigorously prove that if gradient descent is adopted, then ATM iterations implicitly optimize the upper bound of a hypothetical multi-task model trained jointly on the datasets of all tasks.
To summarize, our contribution is four-fold:
  1. (i)
    We show that, under specific conditions, task vectors are equivalent to or approximate the gradients of the corresponding task losses.
  2. (ii)
    We introduce Alternating Tuning and Merging (ATM), a generic framework that generalizes task arithmetic and facilitates more gradual multi-task knowledge integration. Its flexibility allows it to incorporate any conflict-resolution method for further gains.
  3. (iii)
    We propose two ways of applying ATM in practice depending on the constraints. The first (PA-ATM) replaces multi-task learning where data sharing is forbidden, whereas the second (PH-ATM) is a post-hoc enhancement of any multi-task model.
  4. (iv)
    We mathematically prove that throughout the execution of ATM, the evolving base model reduces the loss of a multi-task model finetuned on all tasks jointly.
2 Task Vectors as Gradients
In this section, we formally define the relationship between task vectors and task-loss gradients.
Theorem 1.
Let $\left\{\theta^{(k)}_{t}\right\}_{t=1}^{|T|}$ be a set of models obtained by finetuning the base model $\theta_{\text{base}}$ for $k$ epochs on tasks $T$ using GD with a learning rate $\eta$, where finetuning task $t\in T$ minimizes the loss $\overline{L}_{t}(\theta)=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\ell(x_{i},y_{i},\theta)$. Additionally, let $\left\{\tau_{t}^{(k)}\right\}_{t=1}^{|T|}$ denote the corresponding set of task vectors, with each $\tau^{(k)}_{t}=\theta^{(k)}_{t}-\theta_{\text{base}}$. Let $\tau^{(k)}_{\text{MT}}$ be the multi-task vector $\tau^{(k)}_{\text{MT}}=\sum_{t\in T}\tau_{t}^{(k)}$. Finally, let $\theta_{\text{MT}}^{(k)}$ represent the model obtained by minimizing the combined loss $\sum_{i=1}^{|T|}\overline{L}_{i}$ for $k$ epochs using GD with a learning rate of $\alpha\eta$. It holds that
$\tau^{(1)}_{\text{MT}}=-\eta\nabla\sum_{t\in T}\overline{L}_{t}(\theta_{\text{base}})$ (1)
$\tau^{(k)}_{\text{MT}}=-\eta\sum_{t\in T}\sum_{j=0}^{k-1}\nabla\overline{L}_{i}(\theta_{\text{MT}}^{(j)})+\frac{\eta^{2}}{2}C(\{\theta_{\text{MT}}^{(j)}\}_{j=1}^{k-2})+O(\eta^{3})$ (2)
with
$C(\{\theta_{\text{MT}}^{(j)}\}_{j=1}^{h})=\sum_{t\in T}\sum_{\ell=0}^{h}\nabla^{2}\overline{L}_{t}(\theta^{(\ell)}_{\text{MT}})\sum_{m=0}^{\ell}\left[\alpha\sum_{t^{\prime}\neq t,t^{\prime}\in T}\nabla\overline{L}_{t}^{\prime}(\theta^{(m)}_{\text{MT}})+(\alpha-1)\nabla\overline{L}_{t}(\theta^{(m)}_{\text{MT}})\right]$ (3)
We provide the proof in Appendix A.1. To better appreciate the relationship between a task vector and the gradient computed on the corresponding task dataset, consider the single-task case, where the task vector is exactly the additive inverse of the gradient, scaled by the learning rate $\eta$.
Figure 2: Task arithmetic accuracy: $1$ epoch vs.convergence finetuning
Remark 1.
From Theorem 1, it follows that, for a single task $t$, and after a single finetuning epoch,
$\tau_{t}=-\eta\nabla\overline{L}_{t}(\theta_{\text{base}})$ (4)
where $\eta$ is the learning rate.
This implies that, under the theorem’s assumptions, adding the task vector to the pretrained model approximates the effect of finetuning the model; the same intuition applies to the multi-task case.
In fact, in Fig. 2 we report evidence that the multi-task model obtained by merging models finetuned for a single epoch performs competitively to the one obtained by merging models finetuned to convergence across all tasks. In other words, for the sake of multi-task model merging, performing one epoch of finetuning is often enough, being the task vectors better approximations of the gradients. Further discussion can be found in appendix 9.
We formalize this through the following corollary.
Corollary.
Let $\theta_{\text{TA}}^{(k)}=\theta_{\text{base}}+{\alpha}\sum_{t=1}^{T}\tau_{t}^{(k)}$ be the model obtained using vanilla task arithmetics. Using the same notation of Theorem 1, it holds that
$\theta_{\text{TA}}^{(1)}=\theta_{\text{MT}}^{(1)}$ (5)
$\theta_{\text{TA}}^{(k)}=\theta_{\text{MT}}^{(k)}+\frac{\eta^{2}}{2}C(\{\theta_{\text{MT}}^{(j)}\}_{j=1}^{k-2})+O(\eta^{3})\;\;\;\text{for}\ k>1$ (6)
The corollary indicates that for $k>1$, the multi-task vector still approximates the gradient of a model finetuned for the same number of epochs via GD, with an error on the order of $o(\eta^{2})$. Despite the lack of exact equivalence, the effectiveness of task vectors still holds and can be explained by the fact that most of the model’s finetuning trajectory is driven by the gradient from the first epoch.
In Fig. 3(a) we plot the epoch-wise normalized gradient norm $\frac{\|\nabla_{\theta}^{(k)}L\|}{\sum_{k^{\prime}=1}^{K}\|\nabla_{\theta}^{(k^{\prime})}L\|}$. We observe that the first epoch contributes the most, accounting for up to $70\%$ of the total gradient norms across all epochs. In cases where this is not true, such as with the RESISC45 dataset, we speculate that the direction is still largely determined by the first epoch. As depicted in Fig. 3(b), the gradients from the first $5$ epochs maintain a high cosine similarity $(>0.8)$ with the first epoch’s gradient.
It is worth noting that this analysis does not exactly apply when optimizers like Adam are used instead of GD. However, treating Adam as an approximation to GD preserves intuition. From this view, reducing task interference corresponds to minimizing gradient conflicts in multi-task learning.
(a) Normalized gradient norms after $5$ epochs of finetuning.
(b) Pairwise cosine similarities of the gradients of the first $10$ epochs over the RESISC45 dataset.
Figure 3: Analysis of gradient norms and alignment.
3 ATM: Alternating Tuning and Merging
Building upon the insights of Section 2, we argue that task arithmetic is an approximation to a single GD step over the union of all the tasks. Following this parallel, we advocate taking further update steps iteratively.
The overall framework of ATM is depicted in Fig. 1. Specifically, we start from an initial base model $\theta_{\text{base}}^{(0)}$, we finetune it separately on each task to obtain the first-iteration task vectors $\tau_{1}^{(1)},\dots,\tau_{|T|}^{(1)}$. These are then aggregated and added to the base model to form the next-iteration unified model $\theta_{\text{base}}^{(1)}$. The procedure is iterated according to the following equation:
$\theta_{\text{base}}^{(k+1)}=\theta_{\text{base}}^{(k)}+\frac{\alpha}{|T|}\sum_{t\in T}\tau_{t}^{(k)}\quad\forall k=0,\dots,{K-1}\,.$ (7)
The $k$-th iteration task vector for task $t$ is obtained as $\tau^{(k)}_{t}=\theta_{t}^{(k)}-\theta_{\text{base}}^{(k)}$, where $\theta_{t}^{(k)}$ is a model obtained finetuning the $k$-th iteration base model $\theta_{\text{base}}^{(k)}$ on task $t$. The total number of iterations $K$ can be predefined or based on a stop condition.
In practice, each iteration of ATM involves finetuning the current base model on all $|T|$ tasks using their task-specific data, thereby obtaining $|T|$ task vectors. These task vectors approximate the true task-specific directions the current base model should follow in order to attain enhanced performance on the corresponding tasks.
ATM’s merging step at a given iteration consists of adding the mean of current-iteration task vectors to the current base model, although its flexibility allows any interference-resolution method in the task vector literature to be integrated. This step is intended to pull the base model closer to the multi-task basin on the loss landscape. Taking the average over $T$ ensures that the magnitude of the update remains insensitive to the number of tasks. Note that the task vectors from previous iterations can be safely discarded after each iteration. Therefore, at any iteration, ATM only stores the current base model and one task vector for each of the $|T|$ tasks, incurring no additional storage requirements compared to task arithmetic.
ATM is generic and can be applied in numerous ways. One way is to build a multi-task model from scratch using non-centralized training data (PA-ATM) and a pretrained initialization. This tightly relates to FedAVG (McMahan et al., 2017) in the scenario of federated learning. The other way is to improve an existing merged model with a small set of validation data (PH-ATM). The difference between the two lies in the model initialization $\theta_{\text{base}}^{(0)}$ and data usage.
3.1 PA-ATM Upper bounds the Multi-task Loss
We now show that a single PA-ATM step under GD minimizes the mean of average losses ($\bar{L}_{i}$ ) across tasks and prove that, under certain conditions, reducing this objective also decreases the loss of a model trained jointly on all datasets. As in Section 2, we assume GD is used for optimizing the model parameters. This assumption removes the stochasticity of the random sampling, enabling a more straightforward analysis while still providing insights into the dynamics of the optimization process. This section provides a concise overview; the detailed derivation is given in Appendix A.2.
We denote with $t$ both the task and its corresponding dataset, whose cardinality is $n_{t}$. The total number of samples for all tasks is given by $N=\sum_{t\in T}n_{t}$. Inspired by Daheim et al. (2023), we define the target loss as the loss of a model trained jointly on all the datasets: $L_{\text{target}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\ell(x_{i},y_{i},\theta)$.
As detailed in Appendix A.2, when merging occurs after one step of finetuning on each dataset, the PA-ATM update corresponds to performing a GD step over the loss $L_{PA-ATM}=\frac{1}{|T|}\sum_{t\in T}\overline{L_{t}}$.
Having established that one step of PA-ATM in GD minimizes $L_{\text{ATM}}$ a key question follows: under what conditions does minimizing $L_{\text{PA-ATM}}$ also lead to a reduction in $L_{\text{target}}$?
We show that if the drop in the PA-ATM loss exceeds a threshold, dependent on the sizes of the largest and smallest datasets whose losses respectively decrease and increase, then $L_{\text{target}}$ also decreases. In practice, this typically happens whenever the largest dataset’s loss decreases. A rigorous statement of the theorem and its derivation can be found in Appendix A.2.
4 Experiments
Figure 4: PA-ATM vs Baselines as computational budget $K$ varies.
4.1 Experimental setting
Datasets and Model
The backbone model in all our experiments is the ViT-B-16 (Dosovitskiy et al., 2021). We adhere to Ilharco et al. (2022) and compare PA-ATM and PH-ATM against the baselines on the the following image classification datasets: DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), GTSRB (Houben et al., 2013), MNIST (Lecun et al., 1998), RESISC45 (Cheng et al., 2017), SUN397 (Xiao et al., 2010), and SVHN (Netzer et al., 2011).
Baselines and Setup
To gauge the performance of PA-ATM and PH-ATM, we consider several model merging baselines, including task arithmetic (TA) (Ilharco et al., 2022), TIES-merging (Yadav et al., 2023), model breadcrumbs (Davari & Belilovsky, 2023), and DARE (Yu et al., 2023). To ensure fair comparison across experiments, we adhere to the author-recommended hyperparameters whenever needed or available. Specifically, we adhere to Ilharco et al. (2022) and use AdamW (Loshchilov & Hutter, 2019) as the optimizer across all frameworks, including ours. For TIES-merging, we retain the top $15\%$ of weights based on magnitude ranking. For model breadcrumbs, we set $\beta=0.85$ and $\gamma=0.993$. For DARE merging, we use a drop rate of $0.9$. In all settings, we adopt mean aggregation of task vectors and use a scaling factor of $1$ when applying them to the base model. For PA-ATM, we initialize the base model as the ViT-B-16 pretrained on ImageNet (Deng et al., 2009), whereas for PH-ATM, we initialize the base model as the multi-task model obtained from the application of task arithmetic. Note that while the goal of PA-ATM is to build a multi-task model from scratch, the goal of PH-ATM is to improve upon task arithmetic starting in the same conditions as the baseline conflict-resolution methods.
Strategy
Following Ilharco et al. (2022), we evaluate all methods on the held-out test sets provided by each individual dataset and build the validation set by sampling $10\%$ of the training set. We adopt classification accuracy as the evaluation metric, which is standard in the literature.
4.2 Impact of Epoch Distribution on Performance
Figure 5: PA-ATM multi-task accuracy across budget distributions.
Throughout the paper, we adhered to the convention of merging after just a single epoch of finetuning during ATM iterations. One might question whether more epochs per merging step is beneficial. To show that the one-epoch convention is indeed optimal, we conduct the following experiment with PA-ATM. We establish a fixed compute budget of $10$ finetuning epochs for each task, we then seek the optimal distribution of epochs among different numbers of ATM iterations. To exemplify, if $10$ epochs are distributed among $5$ iterations, then in each iteration a task is finetuned for $2$ epochs.
As depicted in Fig. 5, with a fixed compute budget, maximizing iterations while minimizing epochs per iteration yields the best results for ATM. This indicates that more gradual updates to the base model are preferable to abrupt ones. We observe a clear monotonous increase as more iterations are performed. Splitting $10$ epochs across $10$ iterations achieves the highest average accuracy of $88\%$, outperforming the $1$ iteration of $10$ epochs setting (analogous to task arithmetic) by around $16\%$. Detailed results can be found in the appendix 3. Therefore, the one-epoch convention holds and we subsequently adopt it by default.
4.3 Effect of Finetuning Epochs
Figure 6: Average multi-task accuracy as budget varies.
We compare PA-ATM against baselines for different budgets of finetuning epochs. Specifically, we vary the total per-task finetuning epochs ($K$) within $2$, $4$, and $10$ and compare the average test accuracy across tasks. As shown in Fig. 4, PA-ATM consistently outperforms the baselines for all budgets. Interestingly, while PA-ATM’s accuracy improves with more finetuning epochs, the baseline methods remain unaffected; see Fig. 6. In other words, more specialization of the task-specific models does not necessarily benefit the one-shot unified multi-task model. Detailed comparisons can be found in the appendix 2.
[…]
5
4.4 Comparisons in Original Settings
Figure 7: PH-ATM vs Baselines
We compare the PA-ATM and PH-ATM to the baselines under their original data and hyperparameter setups, including author-specified epochs, weight decay, and learning rate warmup. With an abuse of notation for the sake of readability, the epochs for PH-ATM is based on validation data, unlike all other methods that use the training data. As illustrated in Table 4.4, PA-ATM unsurprisingly and significantly outperforms the baselines. With 10 iterations of ATM, it outperforms the best-performing baseline by $15\%$ on average. If the budget is elevated to 30 iterations, the advantage grows to $17\%$.
We emphasize again that all baselines do make use of the validation data for hyperparameter tuning, while PH-ATM leverages it differently. Fig. 7 and Table 4.4 show that PH-ATM also outperforms the baselines across all tasks but one, with an average improvement of $7\%$ over the best baseline. Thus, we postulate that estimating the task-proficiency directions through validation data, as done in PH-ATM, is a better option when further finetuning is allowed.
Method Cars DTD EuroSAT GTSRB MNIST RESISC45 SVHN SUN397 Average
Pretrained 0.64 0.45 0.54 0.43 0.51 0.65 0.51 0.65 0.55
Finetuned 0.87 0.98 0.98 0.99 0.99 0.96 0.97 0.78 0.94
[…]
TIES 0.88 0.45 0.44 0.41 0.63 0.63 0.42 0.63 0.56
Breadcrumbs 0.69 0.52 0.83 0.61 0.95 0.75 0.74 0.68 0.72
DARE 0.69 0.52 0.83 0.61 0.95 0.75 0.76 0.68 0.72
[…]
PA-ATM10 0.79 0.69 0.98 0.97 0.99 0.90 0.96 0.75 0.88
PA-ATM30 0.82 0.76 0.99 0.99 0.99 0.94 0.97 0.76 0.90
Table 1: Accuracy comparison under original baseline settings. PA-ATM10 refers to PA-ATM employing $10$ iterations.
5 Discussion
5.1 Educated Trajectory
Figure 8: PCA of checkpoints
All baselines perform the aggregation step abruptly in an one-shot fashion over the initial pretrained checkpoint, likely overshooting the multi-task optimum. ATM variants, however, gradually traverses the loss landscape as the base model evolves, leading to more informed nudges toward the multi-task optimum. Fig. 8 depicts the 2D PCA projection of various merged checkpoints. Notably, TIES and Model Breadcrumbs, being deterministic post-hoc enhancement of task arithmetic, end up around the same suboptimal basin. DARE, on the other hand, is located far from the rest due to its stochastic pruning. PA-ATM and PH-ATM, however, start from different initializations and take gradual steps toward different and better basins, signaling the effectiveness of iterative merging.
5.2 Task Proficiency is not Mergeability
Figures 6 and 2 show that task-specific expertise does not guarantee better multi-task performance. Stronger task-specific models attain higher task-specific specialization at the cost of deviating more from the gradient approximation; hence, the downstream performance is not a predictor of post-merging success. We hypothesize that highly specialized models diverge in parameter space, making abrupt one-shot merging suboptimal (Fig. 8). From a gradient perspective, this degradation occurs because vanilla task vectors approximate increasingly noisy multi-task gradients as finetuning progresses. A lower degree of specialization keeps models closer to the pretrained checkpoint, as observed by Ortiz-Jimenez et al. (2024) in the tangent space, leading to less aggressive and more mergeable updates. This aligns with Lu et al. (2024), who attribute the issue to vanishing common knowledge in specialized models. Gradient approximation theory further supports this: fewer finetuning epochs yield better multi-task gradient approximations. ATM capitalizes on this by gradually integrating task-specific knowledge, ensuring less aggressive yet iterative merging. At each step, ATM task vectors provide adaptive nudges to the base model without reverting to the original pretrained checkpoint, unlike baseline methods.
5.3 Limitations of Iterative Merging
Although ATM offers strong performance, its iterative nature exhibits drawbacks. It does not immediately provide task representations, as it iteratively generates vectors per task. However, it remains the best option for obtaining the best-performing multi-task model. We ensured a fair comparison by using the same finetuning epochs and data requirements and showed that a small validation set suffices. Lastly, ATM preserves a key advantage of task arithmetic: by generating task vectors independently, it avoids centralizing data, making it suitable for privacy-sensitive settings.
6 Related work
Mode Connectivity and Model Merging
Mode connectivity explores how weights characterize local minima in the loss landscape. Frankle et al. (2020) studied linear mode connectivity in models with shared initialization, while Entezari et al. (2022) showed that models converge to a shared basin after resolving neuron permutations. Permutation-based merging leverages this insight to combine models without ensembling overhead. Singh & Jaggi (2020) proposed optimal-transport weight matching, Git Re-Basin(Ainsworth et al., 2022) introduced three rebasin variants, and REPAIR(Jordan et al., 2023) reduced barriers via activation renormalization. More recently, Navon et al. (2023) explored embedding-space merging, while Crisostomi et al. (2024) introduced cycle-consistent matching. For models with shared initialization, Wortsman et al. (2022) used simple averaging, while Jolicoeur-Martineau et al. (2023) aligned models to the population mean for stability. Weighted averaging approaches, such as RegMean (Jin et al., 2022) and Fisher-weighted averaging (Matena & Raffel, 2021), optimize merging weights based on specific criteria. Daheim et al. (2023) linked gradient mismatch to post-merging multi-task performance. Lastly, Choshen et al. (2022) proposed merging as an alternative to pretraining, arguing that a merged model can outperform any single fine-tuned checkpoint.
Task Vectors
Task vector-based merging (Ilharco et al., 2022) finetunes a pretrained model on different tasks to obtain task vectors (differences between finetuned and original checkpoints). Arithmetic operations on these vectors enable forgetting, analogy learning, and multi-task learning. Several works aim to improve task vector merging by reducing task interference (Deep et al., 2024; Wang et al., 2024; Huang et al., 2024). Some methods include sparsifying task vectors or finetuning only lottery tickets (Panda et al., 2024). TIES-merging (Yadav et al., 2023) merges vectors by pruning, selecting a unified sign vector, and merging disjointly, while Model Breadcrumbs (Davari & Belilovsky, 2023) prunes both small and large-magnitude weights. DARE Merging (Yu et al., 2023) randomly masks out a portion of weights and scales up the rest. AdaMerging (Yang et al., 2023) optimizes aggregation coefficients, while Yang et al. (2024) proposed task-specific modules for test-time adaptation. Ortiz-Jimenez et al. (2024) introduced the concept of weight disentanglement and recommended finetuning in the tangent space. Unlike these one-shot methods, we propose an iterative approach that progressively refines the base model for better multi-task performance.
Federated Averaging
PA-ATM is tightly related to Federated Averaging (FedAVG)(McMahan et al., 2017), as both alternate between local updates and aggregation while starting from a shared initialization. However, their underlying perspective differs. FedAVG is designed for distributed training on private, non-IID data, whereas PA-ATM focuses on model merging by iteratively integrating task-specific updates. In a sense, PA-ATM represents a bridge between model merging and federated learning which originate from two different purposes.
7 Conclusions
This paper connects task vectors to gradients in multi-task learning, forming the foundation for our proposed Alternating Tuning and Merging (ATM) framework. ATM iteratively refines model merging, addressing the shortcomings of one-shot methods. We introduce two practical frameworks: PA-ATM, which generalizes FedAVG and bridges model merging and federated learning, serving primarily for analysis, and PH-ATM, which enhances existing multi-task models using validation data only. Experiments on vision benchmarks confirm that ATM achieves strong performance with efficiency comparable to baselines. Our theoretical analysis further shows that PA-ATM optimizes an upper bound on the multi-task loss, providing insights into ATM’s effectiveness. The flexibility of ATM allows the additional integration of conflict-resolution and gradient-based techniques.
Ethics Statement
This research was conducted with a strong commitment to ethical standards in both data usage and experimental methodology. All datasets utilized in this study are publicly available. No personally identifiable information was accessed or used during the course of this research. Additionally, the experiments were designed to ensure fair comparisons across methods. We encourage future work that adheres to these same principles and addresses the broader societal impacts of machine learning.
Reproducibility Statement
We are committed to ensuring the reproducibility of our results and have taken steps to facilitate this for the broader research community. The code, datasets, and configurations used for the experiments in this paper are made available via a public repository. We are open to providing further instructions on the usage of our code. The hyperparameters, frameworks, and evaluation metrics have been thoroughly documented, and we provide clear descriptions of our experimental setup to allow for straightforward replication of our findings. We encourage the community to utilize these resources and provide feedback for further improvements.
Acknowledgments
This work is supported by the ERC grant no.802554 (SPECGEO), PRIN 2020 project no.2020TA3K9N (LEGO.AI), and PNRR MUR project PE0000013-FAIR.
Appendix A Proofs
A.1 Proofs of Theorem 1 and Corollary Corollary
In this section, we provide proofs for Theorem 1 and Corollary Corollary. For clarity, we restate both the theorem and corollary.
Theorem.
Let $\left\{\theta^{(k)}_{t}\right\}_{t=1}^{|T|}$ be a set of models obtained by finetuning the base model $\theta_{\text{base}}$ for $k$ epochs on tasks $T$ using GD with a learning rate $\eta$, where finetuning task $t\in T$ minimizes the loss $\overline{L}_{t}(\theta)=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\ell(x_{i},y_{i},\theta)$. Additionally, let $\left\{\tau_{t}^{(k)}\right\}_{t=1}^{|T|}$ denote the corresponding set of task vectors, with each $\tau^{(k)}_{t}=\theta^{(k)}_{t}-\theta_{\text{base}}$. Let $\tau^{(k)}_{\text{MT}}$ be the multi-task vector $\tau^{(k)}_{\text{MT}}=\sum_{t\in T}\tau_{t}^{(k)}$. Finally, let $\theta_{\text{MT}}^{(k)}$ represent the model obtained by minimizing the combined loss $\sum_{i=1}^{|T|}\overline{L}_{i}$ for $k$ epochs using GD with a learning rate of $\alpha\eta$. It holds that
$\tau^{(1)}_{\text{MT}}=-\eta\nabla\sum_{t\in T}\overline{L}_{t}(\theta_{\text{base}})$ (8)
$\tau^{(k)}_{\text{MT}}=-\eta\sum_{t\in T}\sum_{j=0}^{k-1}\nabla\overline{L}_{i}(\theta_{\text{MT}}^{(j)})+\frac{\eta^{2}}{2}C(\{\theta_{\text{MT}}^{(j)}\}_{j=1}^{k-2})+O(\eta^{3})$ (9)
with
$C(\{\theta_{\text{MT}}^{(j)}\}_{j=1}^{h})=\sum_{t\in T}\sum_{\ell=0}^{h}\nabla^{2}\overline{L}_{t}(\theta^{(\ell)}_{\text{MT}})\sum_{m=0}^{\ell}\left[\alpha\sum_{t^{\prime}\neq t,t^{\prime}\in T}\nabla\overline{L}_{t}^{\prime}(\theta^{(m)}_{\text{MT}})+(\alpha-1)\nabla\overline{L}_{t}(\theta^{(m)}_{\text{MT}})\right]$ (10)
Corollary.
Let $\theta_{\text{TA}}^{(k)}=\theta_{\text{base}}+{\alpha}\sum_{t=1}^{T}\tau_{t}^{(k)}$ be the model obtained using vanilla task arithmetics. Using the same notation of Theorem 1, it holds that
$\theta_{\text{TA}}^{(1)}=\theta_{\text{MT}}^{(1)}$ (11)
$\theta_{\text{TA}}^{(k)}=\theta_{\text{MT}}^{(k)}+\frac{\eta^{2}}{2}C(\{\theta_{\text{MT}}^{(j)}\}_{j=1}^{k-2})+O(\eta^{3})\;\;\;\text{for}\ k>1$ (12)
We recall that $\theta^{(k)}_{i}$ is the model obtained by finetuning on task $i$ for $k$ epochs, and that both the finetuing on different tasks and the training on the average loss start from a pretrained model $\theta_{\text{base}}$.
To prove the statement of the theorem and of the corollary we need a intermediate result. We introduce the following notation:
$r_{i}(\theta)=\alpha\sum_{j\neq i}\nabla\overline{L}_{j}(\theta)+(\alpha-1)\nabla\overline{L}_{i}(\theta)=\alpha\sum_{j=1}^{|T|}\nabla\overline{L}_{j}(\theta)-\nabla\overline{L}_{i}(\theta_{\text{base}})$ (13)
$p_{i}^{k}(\theta_{\text{base}},\theta^{(1)}_{\text{MT}},\dots,\theta^{(k)}_{\text{MT}})=\sum_{j=0}^{k}r_{i}(\theta^{(k)}_{\text{MT}})$ (14)
$s^{k}_{i}(\theta_{\text{base}},\dots,\theta^{(k)}_{\text{MT}})=\sum_{j=0}^{k}\nabla^{2}\overline{L}_{i}(\theta^{(j)}_{\text{MT}})[p^{j}_{i}(\theta_{\text{base}},\dots,\theta^{(j-1)}_{\text{MT}})].$ (15)
Lemma.
Using the notation introduced in Theorem 1, it holds that
$\theta^{(1)}_{i}=\theta^{(1)}_{\text{MT}}+\eta p_{i}^{0}(\theta_{\text{base}})$ (16)
and for $m\geq 2$
$\theta^{(m+1)}_{i}=\theta^{(m+1)}_{\text{MT}}+\eta p_{i}^{m}(\theta_{\text{base}},\dots,\theta^{(m)}_{\text{MT}})-\frac{\eta^{2}}{2}s^{m-1}_{i}(\theta_{\text{base}},\dots,\theta^{(m-1)}_{\text{MT}})+O(\eta^{3})$ (17)
Proof.
We first show that the statement is true for $m=1$, and then prove the results for $m\geq 2$ by induction. In this case, the base case is given for $m=2$. In the induction step, instead, we prove that if the statement holds for any given case $m$ then it must also hold for the next case $m+1$.
$m=1$. First epoch
For each task $i=1,\dots,|T|$
$\theta^{(1)}_{i}=\theta_{\text{base}}-\eta\nabla\overline{L}_{i}(\theta_{\text{base}})\text{ while }\theta^{(1)}_{\text{MT}}=\theta_{\text{base}}-\alpha\eta\sum_{i\in T}\nabla\overline{L}_{i}(\theta_{\text{base}}).$
Consequently, it holds that
$\theta^{1}_{i}$ $=\theta^{(1)}_{\text{MT}}+\eta\left[\alpha\sum_{j\neq i}\nabla\overline{L}_{j}(\theta_{\text{base}})+(\alpha-1)\nabla\overline{L}_{i}(\theta_{\text{base}})\right]$
$=\theta^{(1)}_{\text{MT}}+\eta r_{i}(\theta_{\text{base}})=\theta^{(1)}_{\text{MT}}+\eta p_{i}^{0}(\theta_{\text{base}}).$
$m=2$. Second epoch
$\theta^{(2)}_{i}$ $=\theta^{(1)}_{i}-\eta\nabla\overline{L}_{i}(\theta^{(1)}_{i})$
$=\theta^{(1)}_{\text{MT}}+{\eta}r_{i}(\theta_{\text{base}})-\eta\nabla\overline{L}_{i}\left(\theta^{(1)}_{\text{MT}}+{\eta}r_{i}(\theta_{\text{base}})\right)$
$\overset{Taylor}{\approx}\theta^{(1)}_{\text{MT}}+{\eta}r_{i}(\theta_{\text{base}})-\eta\nabla\overline{L}_{i}(\theta^{(1)}_{\text{MT}})-\frac{\eta^{2}}{2}\nabla^{2}\overline{L}_{i}(\theta_{\text{MT}}^{(1)})r_{i}(\theta_{\text{base}})+O(\eta^{3})$
$=\theta^{(1)}_{\text{MT}}-\eta\nabla\overline{L}_{i}(\theta^{(1)}_{\text{MT}})+\eta r_{i}(\theta_{\text{base}})-\frac{\eta^{2}}{2}\nabla^{2}\overline{L}_{i}(\theta_{\text{MT}}^{(1)})r_{i}(\theta_{\text{base}})+O(\eta^{3})$
$=\theta^{(1)}_{\text{MT}}-\eta\nabla\overline{L}_{i}(\theta^{(1)}_{\text{MT}})+\eta\alpha\sum_{t\in T}\nabla\overline{L}_{i}(\theta^{(1)}_{\text{MT}})-\eta\alpha\sum_{t\in T}\nabla\overline{L}_{i}(\theta^{(1)}_{\text{MT}})+\eta r_{i}(\theta_{\text{base}})$
$-\frac{\eta^{2}}{2}\nabla^{2}\overline{L}_{i}(\theta_{\text{MT}}^{(1)})r_{i}(\theta_{\text{base}})+O(\eta^{3})$
$=\theta^{(1)}_{\text{MT}}+\eta r_{i}(\theta^{(1)}_{\text{MT}})+\eta r_{i}(\theta_{\text{base}})-\frac{\eta^{2}}{2}\nabla^{2}\overline{L}_{i}(\theta_{\text{MT}}^{(1)})r_{i}(\theta_{\text{base}})+O(\eta^{3})$
$=\theta^{1}_{\text{MT}}+\eta p_{i}^{1}(\theta_{\text{base}},\dots,\theta^{(1)}_{\text{MT}})-\frac{\eta^{2}}{2}s^{0}_{i}(\theta_{\text{base}})+O(\eta^{3})$
Inductive step
Let us assume that
$\theta_{i}^{(m)}$ $=\theta^{(m)}_{\text{MT}}+\eta p_{i}^{m-1}(\theta_{\text{base}},\dots,\theta^{(m-1)}_{\text{MT}})-\frac{\eta^{2}}{2}s^{m-2}_{i}(\theta_{\text{base}},\dots,\theta^{(m-2)}_{\text{MT}})+O(\eta^{3})$
We can derive that
$\theta_{i}^{(m+1)}$ $=\theta^{(m)}_{i}-\eta\nabla\overline{L}_{i}(\theta^{(m)}_{i})$
$=\theta^{(m)}_{\text{MT}}+\eta p_{i}^{m-1}(\theta_{\text{base}},\dots,\theta^{(m-1)}_{\text{MT}})-\frac{\eta^{2}}{2}s^{m-2}_{i}(\theta_{\text{base}},\dots,\theta^{(m-2)}_{\text{MT}})-\eta\nabla\overline{L}_{i}(\theta^{m}_{i})+O(\eta^{3})$
$=\theta^{(m)}_{\text{MT}}+\eta p_{i}^{m-1}(\theta_{\text{base}},\dots,\theta^{(m-1)}_{\text{MT}})-\frac{\eta^{2}}{2}s^{m-2}_{i}(\theta_{\text{base}},\dots,\theta^{(m-2)}_{\text{MT}})$
$-\eta\nabla\overline{L}_{i}\left(\theta^{(m)}_{\text{MT}}+\eta p_{i}^{m-1}(\theta_{\text{base}},\dots,\theta^{(m-1)}_{\text{MT}})-\frac{\eta^{2}}{2}s^{m-2}_{i}(\theta_{\text{base}},\dots,\theta^{(m-2)}_{\text{MT}})\right)+O(\eta^{3})$
$=\theta^{(m)}_{\text{MT}}+\eta p_{i}^{m-1}(\theta_{\text{base}},\dots,\theta^{(m-1)}_{\text{MT}})-\frac{\eta^{2}}{2}s^{m-2}_{i}(\theta_{\text{base}},\dots,\theta^{(m-2)}_{\text{MT}})$
$-\eta\nabla\overline{L}_{i}(\theta^{(m)}_{\text{MT}})-\frac{\eta}{2}\nabla^{2}\overline{L}_{i}(\theta^{(m)}_{\text{MT}})\left(\eta p_{i}^{m-1}(\theta_{\text{base}},\dots,\theta^{(m-1)}_{\text{MT}})-\frac{\eta^{2}}{2}s^{m-2}_{i}(\theta_{\text{base}},\dots,\theta^{(m-2)}_{\text{MT}})\right)+O(\eta^{3})$
$=\theta^{(m)}_{\text{MT}}+\eta p_{i}^{m-1}(\theta_{\text{base}},\dots,\theta^{(m-1)}_{\text{MT}})-\frac{\eta^{2}}{2}s^{m-2}_{i}(\theta_{\text{base}},\dots,\theta^{(m-2)}_{\text{MT}})$
$-\eta\nabla\overline{L}_{i}(\theta^{(m)}_{\text{MT}})-\frac{\eta^{2}}{2}\nabla^{2}\overline{L}_{i}(\theta^{(m)}_{\text{MT}})p_{i}^{m-1}(\theta_{\text{base}},\dots,\theta^{(m-1)}_{\text{MT}})+O(\eta^{3})$
$=\theta^{(m+1)}_{\text{MT}}+\eta p_{i}^{m}(\theta_{\text{base}},\dots,\theta^{(m)}_{\text{MT}})-\frac{\eta^{2}}{2}s^{m-1}_{i}(\theta_{\text{base}},\dots,\theta^{(m-1)}_{\text{MT}})+O(\eta^{3})$
Proof Theorem and Corollary.
For the first epoch
$\theta^{(1)}_{\text{TA}}=\theta_{\text{base}}+\alpha\sum_{i\in T}\tau^{(1)}_{i}=\theta_{\text{base}}-\eta\alpha\sum_{i\in T}\nabla\overline{L}_{i}(\theta_{\text{base}})$
while, choosing $\alpha\eta$ as learning rate for the loss $\sum_{i\in T}\overline{L}_{i}$ :
$\theta^{(1)}_{\text{MT}}=\theta_{\text{base}}-\alpha\eta\sum_{i\in T}\nabla\overline{L}_{i}(\theta_{\text{base}}).$
So $\theta^{(1)}_{\text{MT}}=\theta^{(1)}_{\text{TA}}$.
For $k\geq 2$, notice that
$\theta^{(k)}_{\text{MT}}=\theta_{\text{base}}-\alpha\eta\sum_{j=0}^{k-1}\nabla\sum_{t\in T}\overline{L}_{i}(\theta^{(j)}_{\text{MT}}).$ (18)
Now, using Lemma Lemma, we get:
$-\alpha\eta$ $\sum_{j=0}^{k-1}\nabla\sum_{t\in T}\overline{L}_{t}(\theta^{(j)}_{\text{MT}})+\eta p_{i}^{k-1}(\eta^{0},\dots,\eta^{k-1}_{\text{MT}})$
$=-\alpha\eta\sum_{j=0}^{k-1}\nabla\sum_{t\in T}\overline{L}_{t}(\theta^{(j)}_{\text{MT}})+\sum_{j=0}^{k-1}r_{i}(\theta^{k}_{\text{MT}})$
$=-\alpha\eta\sum_{j=0}^{k-1}\nabla\sum_{t\in T}\overline{L}_{i}(\theta^{(j)}_{\text{MT}})+\sum_{j=0}^{k-1}\alpha\sum_{j\in T}\nabla\overline{L}_{j}(\theta^{(j)}_{\text{MT}})-\nabla\overline{L}_{i}(\theta^{(j)}_{\text{MT}})$
$=-\eta\sum_{j=0}^{k-1}\nabla\overline{L}_{i}(\theta^{(j)}_{\text{MT}})\,.$
Namely:
$\theta_{i}^{(m+1)}$ $=\theta^{(m+1)}_{\text{MT}}+\eta p_{i}^{m}(\theta_{\text{base}},\dots,\theta^{(m)}_{\text{MT}})-\frac{\eta^{2}}{2}s^{m-1}_{i}(\theta_{\text{base}},\dots,\theta^{(m-1)}_{\text{MT}})+O(\eta^{3})$
$=\theta_{\text{base}}-\alpha\eta\sum_{j=0}^{m}\nabla\sum_{t\in T}\overline{L}_{i}(\theta^{(j)}_{\text{MT}})+\eta p_{i}^{m}(\theta_{\text{base}},\dots,\theta^{(m)}_{\text{MT}})-\frac{\eta^{2}}{2}s^{m-1}_{i}(\theta_{\text{base}},\dots,\theta^{(m-1)}_{\text{MT}})+O(\eta^{3})$
$=\theta_{\text{base}}-\eta\sum_{j=0}^{m}\nabla\overline{L}_{i}(\theta^{(j)}_{\text{MT}})-\frac{\eta^{2}}{2}s^{m-1}_{i}(\theta_{\text{base}},\dots,\theta^{(m-1)}_{\text{MT}})+O(\eta^{3})$
we can rewrite the tasks vectors as
$\tau_{i}^{(k)}$ $=\theta^{(k)}_{i}-\theta_{\text{base}}$ (19)
$=-\eta\sum_{j=0}^{k-1}\nabla\overline{L}_{i}(\theta^{(j)}_{\text{MT}})-\frac{\eta^{2}}{2}s^{k-2}_{i}(\theta_{\text{base}},\dots,\theta^{(k-2)}_{\text{MT}})+O(\eta^{3})$ (20)
Consequently the model obtained with TA is
$\theta^{(k)}_{\text{TA}}$ $=\theta_{\text{base}}+\alpha\sum_{i\in T}\tau^{(k)}_{i}$
$=\theta_{\text{base}}-\eta\alpha\sum_{j=0}^{k-1}\sum_{i\in T}\nabla\overline{L}_{i}(\theta^{(j)}_{\text{MT}})-\alpha\sum_{i\in T}\frac{\eta^{2}}{2}s^{k-2}_{i}(\theta_{\text{base}},\dots,\theta^{(k-2)}_{\text{MT}})+O(\eta^{3})$
$=\theta^{(k)}_{\text{MT}}-\alpha\sum_{i\in T}\frac{\eta^{2}}{2}s^{k-2}_{i}(\theta_{\text{base}},\dots,\theta^{(k-2)}_{\text{MT}})+O(\eta^{3})\,.$
A.2 Upper Bouding the Multi-task Loss
In this section, we explore the relationship between the PA-ATM loss, defined as the mean of average losses over all tasks, and the loss of a model trained jointly on all the datasets. We conduct this analysis using GD, rather than SGDs, thus removing stochasticity and enabling a clearer analysis while retaining the key insights. Denoting with $t$ both the task and its corresponding dataset with cardinality $n_{t}$. The total number of samples for all tasks is given by $N=\sum_{t\in T}n_{t}$.
By Theorem 1, when merging occurs after one step of finetuning on each dataset, the PA-ATM update is given by:
$\theta_{\text{base}}^{(k+1)}=\theta_{\text{base}}^{(k)}+\frac{\alpha}{|T|}\sum_{t\in T}\tau_{t}^{(k)}=\theta_{\text{base}}^{(k)}-\alpha\eta\nabla\left(\frac{1}{|T|}\sum_{t\in T}\overline{L_{t}}\right)\,,$ (21)
which corresponds to performing a GD step over the loss $L_{\text{ATM}}=\frac{1}{|T|}\sum_{t\in T}\overline{L_{t}}$.
Having established that one step of PA-ATM in GD minimizes $L_{\text{PA-ATM}}$, a crucial question arises: under what conditions does minimizing $L_{\text{ATM}}$ imply the minimization of $L_{\text{target}}$? In other words, when can we be certain that optimizing the PA-ATM loss will also minimize the loss associated with training jointly on all datasets?
To study if optimizing the PA-ATM loss will also minimize the loss associated with training jointly on all datasets we first note that $L_{\text{PA-ATM}}$ is an unweighted average of the individual dataset losses, while the target loss is a weighted average:
$L_{\text{target}}(\theta)=\frac{\sum_{t\in T}n_{t}\overline{L}_{t}(\theta)}{\sum_{t\in T}n_{t}}\,.$ (22)
We now analyze the parameter update from $\theta^{(k)}$ to $\theta^{(k+1)}$. For both PA-ATM and target methods, we denote the change in loss, $L_{\text{method}}$, as $\Delta L_{\text{method}}=L_{\text{method}}(\theta^{(k)})-L_{\text{method}}(\theta^{(k+1)}).$ In the following theorem, we prove that if the drop in ATM loss exceeds a threshold $\delta$, the target loss will also decrease. The value of $\delta$ depends on the size of the largest dataset with a decreasing loss and the smallest dataset with an increasing loss. In particular, if the former dataset is larger than the latter, a reduction in PA-ATM loss reduces the target loss. In practice, this is ensured when the loss is reduced on the largest dataset.
Theorem 2.
Let $D=\{t\mid\Delta\overline{L}_{t}>0\}$ be the set of datasets where the loss decreases after a parameter update, and $I=\{t\mid\Delta\overline{L}_{t}\leq 0\}$ be the set of datasets where the loss increases or remains unchanged. If the reduction in the PA-ATM loss satisfies $\Delta L_{\text{PA-ATM}}>\delta$, where
$\delta=\frac{1}{|T|}\left(1-\frac{\min_{t\in I}n_{t}}{\max_{t\in D}n_{t}}\right)\sum_{t\in I}|\Delta\overline{L}_{t}|\,,$
then the target loss $L_{\text{target}}$ will also decrease, i.e., $\Delta L_{\text{target}}>0$.
Proof.
[…]
$\sum_{j\in N}|D_{j}|\Delta\overline{L_{j}}=-\sum_{j\in N}|D_{j}||\Delta\overline{L_{j}}|\,.$
Under the hypothesis that $\Delta L_{\text{PA-ATM}}>\delta$, this implies $\sum_{j=1}^{n}\Delta_{j}>n\delta$. We have that $\sum_{j\in P}\Delta\overline{L_{j}}+\sum_{j\in N}\Delta\overline{L_{j}}=\sum_{j\in P}\Delta\overline{L_{j}}-\sum_{j\in N}|\Delta\overline{L_{j}}|>n\delta$, namely $\sum_{j\in P}\Delta\overline{L_{j}}>n\delta+\sum_{j\in N}|\Delta\overline{L_{j}}|$.
We want to prove that $\Delta L_{\text{target}}>0$. Let us now consider $\Delta L_{\text{target}}>0$ iff $\sum_{j=1}^{n}|D_{j}|\Delta_{j}>0$.
$\sum_{j=1}^{n}|D_{j}|\Delta_{j}$ $=\sum_{j\in P}|D_{j}|\Delta\overline{L_{j}}+\sum_{j\in N}|D_{j}|\Delta\overline{L_{j}}$
$=\sum_{j\in P}|D_{j}|\Delta\overline{L_{j}}-\sum_{j\in N}|D_{j}||\Delta\overline{L_{j}}|$
$>\min_{j\in P}|D_{j}|\sum_{j\in P}\Delta\overline{L_{j}}-\max_{j\in N}|D_{j}|\sum_{j\in N}|\Delta\overline{L_{j}}|$
$>\min_{j\in P}|D_{j}|\left[n\delta+\sum_{j\in N}|\Delta\overline{L_{j}}|\right]-\max_{j\in N}|D_{j}|\sum_{j\in N}|\Delta\overline{L_{j}}|$
$=n\min_{j\in P}|D_{j}|\delta+(\min_{j\in P}|D_{j}|-\max_{j\in N}|D_{j}|\sum_{j\in N})|\Delta\overline{L_{j}}|\,.$
The last line of the previous equation is positive by hypothesis, since we assumed $\delta>\frac{1}{n}(1-\frac{\min_{j\in N}|D_{j}|}{\max_{j\in P}|D_{j}|})\sum_{j\in N}|\Delta\overline{L_{j}}|$. ∎
Remark 2.
If we choose the target loss to be the maximum of the average loss across all datasets $L_{\text{target}}=\max_{t\in T}\overline{L}_{t}$, by leveraging the equivalence between the $L_{1}$-norm and the max norm, we obtain the bound $L_{\text{target}}\leq T\cdot L_{\text{PA-ATM}}$.
Appendix B Additional Results
B.1 Full results over varying computational budget
In the main paper, we pictorially illustrated the multi-task accuracy of the baselines and ATM variants, in the form of radar plots. For deeper analysis, here we report the full results of PA-ATM and PH-ATM compared to the baselines, as the computational budget $K$ varies across 2, 4, 6, 8, and 10 epochs. Again, we point out that, unlike PA-ATM and the baselines, PH-ATM is finetuned only with validation data. See Table 2.
$K$ Method Cars DTD EuroSAT GTSRB MNIST RESISC45 SVHN SUN397 Average
2 Task Arithmetic 0.68 0.50 0.81 0.62 0.95 0.74 0.78 0.68 0.72
TIES 0.75 0.46 0.52 0.44 0.57 0.66 0.52 0.66 0.57
Breadcrumbs 0.68 0.50 0.80 0.61 0.95 0.74 0.77 0.68 0.72
DARE 0.68 0.50 0.81 0.62 0.95 0.74 0.78 0.68 0.72
PA-ATM 0.71 0.53 0.90 0.75 0.98 0.78 0.86 0.70 0.78
4 Task Arithmetic 0.69 0.49 0.79 0.63 0.96 0.74 0.78 0.68 0.72
TIES 0.80 0.46 0.53 0.45 0.58 0.66 0.50 0.65 0.58
DARE 0.68 0.49 0.79 0.63 0.96 0.74 0.78 0.68 0.72
Breadcrumbs 0.68 0.49 0.79 0.62 0.96 0.73 0.77 0.68 0.76
PA-ATM 0.74 0.59 0.95 0.90 0.99 0.84 0.93 0.72 0.83
6 Task Arithmetic 0.69 0.49 0.80 0.64 0.96 0.74 0.79 0.68 0.72
TIES 0.82 0.45 0.50 0.45 0.57 0.66 0.51 0.65 0.58
DARE 0.69 0.49 0.80 0.63 0.96 0.74 0.79 0.68 0.72
Breadcrumbs 0.69 0.48 0.79 0.63 0.96 0.73 0.77 0.68 0.72
PA-ATM 0.77 0.64 0.97 0.95 0.99 0.87 0.95 0.74 0.86
8 Task Arithmetic 0.70 0.49 0.79 0.65 0.96 0.73 0.78 0.68 0.72
TIES 0.83 0.46 0.52 0.45 0.60 0.66 0.49 0.65 0.58
DARE 0.69 0.49 0.79 0.64 0.96 0.73 0.79 0.68 0.72
Breadcrumbs 0.69 0.49 0.79 0.64 0.96 0.73 0.77 0.68 0.72
PA-ATM 0.78 0.67 0.98 0.96 0.99 0.89 0.96 0.74 0.87
10 Task Arithmetic 0.69 0.49 0.80 0.64 0.97 0.73 0.79 0.68 0.72
TIES 0.83 0.46 0.52 0.44 0.60 0.65 0.51 0.65 0.58
DARE 0.68 0.49 0.80 0.63 0.97 0.73 0.79 0.68 0.72
Breadcrumbs 0.69 0.49 0.80 0.64 0.97 0.73 0.77 0.68 0.72
PA-ATM 0.79 0.69 0.98 0.97 0.99 0.90 0.96 0.75 0.88
Table 2: ATM vs. Baselines as budget varies (ViT-B-16)
B.2 Full results of varying ATM budget distribution
We show the full accuracy results as we vary the distribution of 10 epochs into different numbers of ATM iterations. See Table 3.
Iterations Epochs/iteration Cars DTD EuroSAT GTSRB MNIST RESISC45 SVHN SUN397 Average
1 10 0.6886 0.4858 0.8027 0.6402 0.9681 0.7276 0.7850 0.6792 0.7222
2 5 0.7203 0.5337 0.9093 0.8004 0.9862 0.7857 0.8971 0.6964 0.7911
5 2 0.7668 0.6126 0.9616 0.9395 0.9921 0.8716 0.9494 0.7279 0.8527
10 1 0.7904 0.6897 0.9837 0.9713 0.9943 0.9032 0.9644 0.7490 0.8808
Table 3: PA-ATM accuracies for different budget distributions
Appendix C Task difficulty related to epochs
To gain further insight into when finetuning for more epochs might offer advantages in Task Arithmetic (TA), we examined correlations between task “difficulty” and the difference in merged performance under two settings.
Specifically, we tracked both loss and accuracy at the first and last epochs during finetuning. For each task, we then derived four key metrics:
acc_gap $=\text{acc}_{\text{last}}-\text{acc}_{\text{first}},$
acc_ratio $=\frac{\text{acc}_{\text{first}}}{\text{acc}_{\text{last}}},$
loss_gap $=\text{loss}_{\text{first}}-\text{loss}_{\text{last}},$
normalized_loss_gap $=\frac{\texttt{loss\_gap}}{\text{loss}_{\text{first}}}.$
Intuitively, tasks showing smaller loss_gap or acc_gap can be viewed as “easier” relative to the pre-trained initialization, whereas larger gaps indicate a need for more substantial finetuning to reach high accuracy.
Next, we sampled approximately 800 distinct five-task subsets drawn from the twenty tasks used in (Gargiulo et al., 2024), merged the individual fine-tuned models using both single-epoch TA (1 epoch) and vanilla TA, and measured the average test accuracy of each merged model.
(a) Loss gap
(b) Accuracy gap
(c) Accuracy ratio
(d) Normalized loss gap
Figure 9: Visualization of task difficulty metrics against delta of normalized merged accuracy.
From Figure 9, some key observations arise. For “easier” tasks, characterized by smaller acc_gap or loss_gap, the difference $\Delta\text{avg}$ is close to zero, indicating that single-epoch TA typically matches or comes very close to vanilla TA. In such settings, the tasks do not heavily diverge from the pretrained representation, and simpler merging strategies appear sufficient. For “harder” tasks, where acc_gap or loss_gap is larger, we observe that $\Delta\text{avg}$ tends to be negative, favoring vanilla TA. Here, specialized finetuning steps seem more crucial, giving an edge to vanilla TA in preserving task-specific refinements within the merged model.
Overall, these findings suggest that while vanilla TA or standard multi-task approaches may be more suitable for especially challenging tasks (i.e. those requiring extensive finetuning to reach high accuracy), single-epoch TA is appealing in budget-critical settings or when tasks are relatively straightforward to learn from a pretrained model. In practice, one can anticipate slightly better results from vanilla TA for tasks that significantly diverge from the pretrained initialization, but benefit from single-epoch TA due to less finetuning.